jaiden123232:

Find −a2 − 3b3 + c2 + 2b3 − c2 if a = 3, b = 2, and c = −3.

mhanifa:

You may substitute values and solve

mhanifa:

c^2 should have positive sign, but you can even cancel them before starting calculations.

mhanifa:

−a2 − 3b3 + c2 + 2b3 − c2 = -a2 - b3, this is the first step \[-3^2 - 2^3 = -9 - 8 = -17\] this is the second step

mhanifa:

@extrinix wrote:
There is no equal to, there is only given one side of the equation.
If no 'equal to symbol', then no equation. You are looking for the value of the expression, and you have to use that symbol until you reveal a result.

mhanifa:

@extrinix wrote:
\(−a^2 − 3b^3 + c^2 + 2b^3 − c^2\) So you need to substitute each of the values in, \(a=3\) \(b=2\) \(c=-3\) \(-3^2-3(2)^3+-3^2+2(2)^3-(-3)^2\) Now we can simplify it, \(-9 -24 -9 +16 -9\) So from this we can then solve it, \(-9 -24 -9 +16 -9 = ?\) \(-33 -9 +16 -9 = ?\) \(-42 +16 -9 = ?\) \(-26 -9 = ?\) \(-35 = ?\) With the substitutions you end up with \(-35\).
\(-3^2-3(2)^3+(-3)^2+2(2)^3-(-3)^2\) makes it right as the third and last terms cancelled

mhanifa:

This is the short way, simplification first, substitution next.

Extrinix:

\(−a^2 − 3b^3 + c^2 + 2b^3 − c^2\) So you need to substitute each of the values in, \(a=3\) \(b=2\) \(c=-3\) \(-3^2-3(2)^3+-3^2+2(2)^3-(-3)^2\) Now we can simplify it, \(-9 -24 +16\) So from this we can then solve it, \(-9 -24 +16 = ?\) \(-33 +16 = ?\) \(-17 = ?\) With the substitutions you end up with \(-17\).

mhanifa:

@extrinix wrote:
Happy?
Thanks, but it is not to make me happy. The person who needs the answer asked and forgot.

jhonyy9:

like a first step add the same terms

jhonyy9:

+c^2 with -c^2 cancel out

jhonyy9:

the second and the 4th term you can adding

jhonyy9:

@jhonyy9 wrote:
+c^2 with -c^2 cancel out
i think this is the first step

jhonyy9:

@jhonyy9 wrote:
the second and the 4th term you can adding
and this the second one

jhonyy9:

first you need simplifie thie equation and just afer that you can substitute these values